OpenStudy (tubber):

I'm lost, please help, will medal & fan (this isn't my first q) Scotty invests an amount at 2% interest compounded annually. After 3 years, he has $3183.62. What was the original amount Scotty invested? $3000 $2750 $2250 $2500

OpenStudy (tubber):

@Awolflover1

OpenStudy (maddielb):

Well first of all, Welcome to OpenStudy!

OpenStudy (tubber):

haha ty I had an account before

OpenStudy (maddielb):

Ah :p

OpenStudy (tubber):

So can you help me solve for original value?

OpenStudy (breanna1999):

The easiest way to figure it out is to plug in the answers to the interest rate formula

OpenStudy (maddielb):

Sorry I knew how to do this when I got here, my brain literally forgot what it was thinking :p Ill let @breanna1999 take it :p

OpenStudy (tubber):

lol that is k ok MaddieLB

OpenStudy (tubber):

So I=pr?

OpenStudy (tubber):

or I=prt?

OpenStudy (welshfella):

A = P(1 + r)^n where A is the amount after n years, P is the Principal ( the original amount) and r is the annual rate.

OpenStudy (tubber):

oh ok

OpenStudy (breanna1999):

so i found that 3000(1.02)^3 = 3183.62

OpenStudy (tubber):

so 3183.62=.... Which is principal again??? =/ ok ty breanna, how do you do that?

OpenStudy (welshfella):

you need plug in the given values and solve for P

OpenStudy (breanna1999):

you have (1 +r) r is the rate which is 0.02

OpenStudy (tubber):

Ok ty guys! Which one of you should I medal...

OpenStudy (welshfella):

3183.62 = P ( 1.02)^3 P = 3183.62 / (1.02^^3

OpenStudy (tubber):

OH ok ty welshfella I think I understand now

OpenStudy (welshfella):

good